OpenStudy (anonymous):

1/4y-3=9 Help me!!plz!!! i got three but the answersheet says its 48

OpenStudy (anonymous):

\[\frac{ 1 }{ 4y-3 }=9\] \[4y-3=\frac{ 1 }{ 9 }\] \[4y=9+\frac{ 1 }{ 3 }=\frac{ 27+1 }{3}=\frac{ 28 }{ 3 },y=\frac{ 28 }{21 }=\frac{ 4 }{ 3 }\]